Angle A and Angle B are complementaey. In addition, angle A is supplementary to angle C. What are the measures of angles A and C is angle B measures 37.

Answer by dabanfield(803)

Angle A and Angle B are complementaey. In addition, angle A is supplementary to angle C. What are the measures of angles A and C is angle B measures 37.

A + B = 90
A + C + 180
So:
If B = 37 then A + 37 = 90.
A = 90 - 37 = 53
If A + C = 180 then
53 + C = 180
C = 180 - 53
C = 127